### v1

# BiTrajDiff: Bidirectional Trajectory Generation with Diffusion Models for Offline Reinforcement Learning

Abstract
Recent advances in offline Reinforcement Learning (RL) have proven that effective policy learning can benefit from imposing conservative constraints on pre-collected datasets. However, such static datasets often exhibit distribution bias, resulting in limited generalizability. To address this limitation, a straightforward solution is data augmentation (DA), which leverages generative models to enrich data distribution. Despite the promising results, current DA techniques focus solely on reconstructing future trajectories from given states, while ignoring the exploration of history transitions that reach them. This single-direction paradigm inevitably hinders the discovery of diverse behavior patterns, especially those leading to critical states that may have yielded high-reward outcomes. In this work, we introduce Bidirectional Trajectory Diffusion (BiTrajDiff), a novel DA framework for offline RL that models both future and history trajectories from any intermediate states. Specifically, we decompose the trajectory generation task into two independent yet complementary diffusion processes: one generating forward trajectories to predict future dynamics, and the other generating backward trajectories to trace essential history transitions. BiTrajDiff can efficiently leverage critical states as anchors to expand into potentially valuable yet underexplored regions of the state space, thereby facilitating dataset diversity. Extensive experiments on the D4RL benchmark suite demonstrate that BiTrajDiff achieves superior performance compared to other advanced DA methods across various offline RL backbones.
1 Introduction
Offline Reinforcement Learning (Offline RL) [11] seeks to derive effective decision policies exclusively from pre-collected datasets, obviating the necessity for online exploration due to the inherent risks and costs associated with real-time interaction [25]. This paradigm has demonstrated notable efficacy across various real-world applications, including robotic control [38, 37, 29] and power grid management [3, 45]. A primary challenge in offline RL is addressing the out-of-distribution (OOD) problem, stemming from the inherent divergence between the static offline dataset distribution and the learned policy behavioral distribution during training. Consequently, the direct application of conventional online RL methods inevitably leads to extrapolation error [25], characterized by the overestimation of value functions for OOD state-action pairs. To tackle the OOD problem, offline RL research emphasizes the development of techniques that constrain learned policies to remain within the distribution of the static offline dataset. These approaches include imposing explicit constraint terms during policy optimization [38, 5] and penalizing the value function estimates for OOD actions [31, 46].
Despite promising advancements in offline RL, current algorithms critically restrict agent actions, as evidenced by [25, 11, 38]. This limitation stems from the inherent mechanism constraining learned policy to the biased distribution of offline datasets, often leading to overfitting on limited behavioral patterns. While this principle effectively facilitates the replication offavorable behavioral patterns within the data, it inevitably curtails generalizability, hindering the exploration of optimal behaviors beyond dataset coverage. A common approach to mitigate the distribution bias is Data Augmentation (DA) method, which enriches the offline dataset with a large amount of synthesized high-fidelity data. For instance, studies like [30, 26, 27] leverage one generative diffusion model to approximate the distribution of single-direction forward-future trajectories for augmenting offline data. Notably, as data augmentation (DA) methods operate by modifying the training dataset, they are orthogonal to specific offline RL algorithms, enabling them to enhance any such algorithm without altering its core optimization mechanisms. However, a critical limitation of DA methods is their exclusive focus on trajectories originating from a predefined initial state. While these generated single-direction trajectories effectively explore behaviors emerging from a given state, they neglect the exploration of trajectories leading to that state. This oversight restricts the utility of the generated trajectory to locally explore the trajectory space immediately surrounding existing data, specifically for transitions originating from states already in the dataset. Consequently, while DA methods successfully generate trajectories for local state pairs already reachable within the offline dataset, they are ineffective in discovering global trajectories for previously unreachable state pairs. This naturally leads us to a neglected yet unresolved question:
Given an intermediate state, can we not only imagine its future trajectory but also reconstruct its possible history trajectory leading to the state?
In this paper, we propose the Bidirectional Trajectory Diffusion framework, abbreviated as BiTrajDiff, a novel data augmentation (DA) approach designed to generate globally diverse behavioral patterns beyond those present in offline datasets. Diverging from conventional DA approaches generating only local, single-direction trajectories starting from states in the dataset, BiTrajDiff synthesizes bidirectional trajectories by concurrent generation of forward-future and backward-history trajectories, conditioned on intermediate anchor states. This paradigm enables global connectivity between originally unreachable states. Technically, we decouple the bidirectional trajectory generation with two diffusion models: a forward trajectory diffusion and a backward trajectory diffusion. The forward model, conditioned on an initial state, generates subsequent state sequences, while the backward model, conditioned on a terminal state, generates preceding state sequences. This design enables us to stitch these two types of generated trajectories by conditioning them on the same intermediate states. To ensure high-quality and complete generated data, a supervised learning-based inverse dynamics model and a reward model are concurrently deployed to imbue the state sequences with action and reward signals, followed by a novel filtering mechanism that excludes OOD and suboptimal trajectories. Our core idea is that synchronously imagining the forward-future and backward-histroy trajectory allows independent exploration of potential transitions from both temporal directions, and their intersection at learned intermediate states creates novel paths that logically “stitch” these separately explored regions, thereby globally connecting states with no direct transitional evidence in the original data. Therefore, augmenting offline RL with BiTrajDiff-generated data can lead to substantial performance improvements.
Our contributions are summarized as follows: (i). We are the first to explore and address the challenge of concurrently imagining both future and historical trajectories on a given state for data augmentation in offline RL. This paradigm enables the acquisition of diverse behavioral patterns. (ii). We propose BiTrajDiff, a novel framework that generates global trajectories between previously unreachable states in the original dataset, achieved by independently generating and then stitching together both forward-future and backward-history trajectories conditioned on common intermediate states. (iii). We conduct extensive experiments on the D4RL benchmark, applying BiTrajDiff across various offline RL algorithms. Our results demonstrate that BiTrajDiff achieves significantly superior performance compared to other advanced DA baselines, highlighting its effectiveness and robustness.
2 Related work
Offline Reinforcement Learning encompasses four primary categories: policy constraint [44, 46], value regularization [23, 24], model-based [48, 2], and return-conditioned supervised learning [4, 34]. Policy constraint methods improve policies by maximizing expected returns while restricting agent actions to the offline dataset, employing techniques such as explicit behavior cloning regularization [38, 5] or implicit optimization towards derived in-sample optimal policies [32, 50]. Meanwhile, value regularization mitigates overestimation of OOD action values by promoting conservative value functions; recent work focuses on establishing precise lower bounds for Q-values to facilitate in-sample action selection [31, 33]. Model-based approaches construct task-oriented world models from offline data, including both one-step dynamics [48, 2] and multi-step sequence models [13, 14], which are subsequently used for offline reward penalization [49, 22] and online decision planning to ensure in-distribution actions [15, 16]. Return-conditioned supervised learning trains policies to generate trajectories conditioned on return values via supervised learning, with transformer [4, 39] and diffusion models [21, 1] serving as common policy backbones to enhance expressiveness and robustness. Nevertheless, a significant limitation of these approaches is their restriction of behavior to the inherent dataset distribution, which inevitably constrains the generalizability of the policies.
Data Augmentation (DA) Method for Offline RL leverages data augmentation to synthesize interaction data for subsequent RL training. To enhance data diversity and precision, Zhang et al. [51] train an ensemble of dynamics models and augment the RL buffer with generated transitions weighted by ensemble variance. Lu et al. [30] first utilize the notable generative capacity of diffusion to flexibly oversample generated one-step transitions for RL training. Subsequently, recent attempts in RL data augmentation have mainly focused on direct forward future trajectory generation. Several studies aim to achieve precise policy improvement by generating on-policy trajectories. For example, Jackson et al. [20] achieve on-policy trajectory generation by computing guidance derived from the policy model, whereas He et al. [17] employ interaction history embeddings to condition the generation process. Beyond on-policy generation, Wang et al. [43] employ diverse criteria, including dynamics error, temporal-difference error, and curiosity, to guide the augmented trajectory generation. Separately, Yang and Wang [47] investigate the generation of reverse trajectories terminated at conditioned states. Furthermore, Li et al. [26] explore stitching similar offline trajectories using intermediate states generated by masked diffusion. Nevertheless, a critical limitation of these DA methods is their inherent reliance on forward-only trajectory generation, which constrains them to local explorations around existing data. In contrast, our BiTrajDiff offers a distinct advantage by concurrently generating forward and backward trajectories starting from an intermediate state.
3 Preliminary
Reinforcement Learning (RL) paradigm [42] is formally defined as a Markov Decision Process (MDP), $\mathcal{M}=\langle\mathcal{S},\mathcal{A},P,r,\gamma,\rho_{0}\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ represents the environment dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in[0,1)$ is the discount factor, and $\rho_{0}$ denotes the initial state distribution. At each timestep $t$, an agent executes an action $a_{t}$ according to its policy $\pi(\cdot|s_{t})$ in the current state $s_{t}$, transitioning to the subsequent state $s_{t+1}$ with reward $r_{t}$ based on $P$. The objective of the agent is to find an optimal policy $\pi^{*}$ that maximizes the expected discounted return: $\pi^{*}=\arg\max_{\pi}\mathcal{J}(\pi)=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$. In contrast to online RL, offline RL [25] constrains the agent from interaction with the environment, relying solely on a static trajectories dataset $\mathcal{D}=\{\tau^{i}\}_{i=1}^{N}$, where $N$ is the number of trajectories and each trajectory $\tau^{i}=\{(s^{i}_{t},a^{i}_{t},r^{i}_{t})\}_{t=0}^{T-1}$ comprises a sequence of state-action-reward tuples.
Diffusion Models [19, 41] are generative frameworks that learn data distributions through a forward diffusion process and a reverse denoising process. The forward process gradually corrupts data $\mathbf{x}_{0}$ via a predefined noise schedule $\{\beta_{k}\}_{k=1}^{K}$, where each transition is defined as $q(\mathbf{x}_{k}|\mathbf{x}_{k-1})=\mathcal{N}\left(\mathbf{x}_{k};\sqrt{1-\beta_{k}}\mathbf{x}_{k-1},\beta_{k}\mathbf{I}\right)$. As $K\to\infty$, the terminal distribution $q(\mathbf{x}_{K})$ converges to an isotropic Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$. The reverse process, parameterized by $\theta$, aims to iteratively reconstruct the original data through leanerable Gaussian transitions: $p_{\theta}(\mathbf{x}_{k-1}|\mathbf{x}_{k})=\mathcal{N}\left(\mathbf{x}_{k-1};\mu_{\theta}(\mathbf{x}_{k},k),\Sigma_{\theta}(\mathbf{x}_{k},k)\right).$ By maximizing the Empirical Lower Bound (ELBO) [40] on the log-likelihood of the sampled data, the diffusion model can be trained with a simplified surrogate loss [40]:
$\mathcal{L}_{\text{denoise}}(\theta)=\mathbb{E}_{\mathbf{x}_{0}\sim q,k\sim\mathcal{U}\{1,K\},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\epsilon_{\theta}(\mathbf{x}_{k},k)-\epsilon\right\|^{2}_{2}\right],$ (1)
where $\mathcal{U}$ denotes the discrete uniform distribution, and $\epsilon_{\theta}$ is the deep model parameterized with $\theta$ to predict the noise. Both the mean $\mu_{\theta}(\mathbf{x}_{k},k)$ and covariance $\Sigma_{\theta}(\mathbf{x}_{k},k)$ of the reverse process are analytically derivable from $\epsilon_{\theta}$. The sampling is then performed by initializing with $\mathbf{x}_{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and iteratively applying the learned reverse transitions.
4 Methodology
This section introduces the BiTrajDiff framework for offline RL, which comprises two core components: bidirectional diffusion training and bidirectional trajectory generation. During the bidirectional diffusion training phase, two distinct diffusion models are employed to model the distributions of forward-future and backward-history state sequences within trajectory data. Conditioned on an intermediate state and cumulative reward signals, these models generate state sequences in corresponding temporal directions. Then, in the subsequent bidirectional trajectory generation phase, the generated forward and backward trajectories are reconciled through filling and filtering operations to produce novel global trajectories that connect previously unreachable states in the original dataset.
4.1 Bidirectional Diffusion Training
Building upon recent advances in diffusion-based sequence generation [8, 12], we propose a bidirectional trajectory generation framework comprising independent forward and backward trajectory diffusion models. These models are separately trained on offline datasets to learn the trajectory distribution of the behavior policy, facilitating bidirectional state trajectory generation from shared intermediate “anchor” states.
Formally, for each diffusion model, the state trajectory is generated simultaneously per time step $t$ over the planning horizon $H$: $\mathbf{x}_{k}(\tau)=\{s_{t},s_{t+1},\cdots,s_{t+H-1}\}_{k}$, where $k$ denotes the denoise timestep and $t$ represents the MDP timestep. In this way, the conditional generative problem via diffusion model can be formulated as:
$\min_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{\tau\sim\mathcal{D}}[\log p_{\theta}(\mathbf{x}_{0}(\tau)|\mathbf{y}(\tau))],$ (2)
where $\mathbf{x}_{0}(\tau)$ denotes the final generated subtrajectory of diffusion, $\mathbf{y}(\tau)$ represents the condition for generation, and $\theta$ represents the learnable parameters of the diffusion model. While various condition metrics are available, such as temporal difference error, Q-function, or cumulative return [43], our framework utilizes the cumulative return $R$ and the state $s$ as conditional input $\mathbf{y}(\tau)$. This approach, consistent with prior works [1, 6], enables the generation of trajectories with desired returns starting from the specific states. However, to enable bidirectional trajectory generation, the formations of the backward and forward diffusion condition inputs are quite distinct.
Specifically, the forward diffusion model generates finite-horizon state trajectories $\mathbf{x}^{f}_{0}(\tau)=\{s_{t},s_{t+1},\cdots,s_{t+H-1}\}$ started at initial state $s_{t}$ with target return $R^{f}(\tau)=\sum_{i=t}^{t+H}\gamma^{i-t}r_{i}$, thus generating physically consistent future trajectories within the offline distribution. Conversely, the backward diffusion model reconstructs historical trajectories $\mathbf{x}^{b}_{0}(\tau)=\{s_{t-H+1},\cdots,s_{t}\}$ that ends on terminal state $s_{t}$ and accrued return $R^{b}(\tau)=\sum_{i=t-H}^{t}\gamma^{t-i}r_{i}$. For brevity and clarity, we utilize superscript-free notation (e.g., $\mathbf{x}_{k}$, $\mathbf{y}$, $R(\tau)$) in the following sections to denote computational processes common to both directional models.
Based on the two distinct conditions, we can model the conditional trajectory generation processes by the corresponding diffusion sampling composed of the forward noising process $q(\mathbf{x}_{k+1}(\tau)|\mathbf{x}_{k}(\tau))$ and the reverse denoising process $p_{\theta}(\mathbf{x}_{k-1}(\tau)|\mathbf{x}_{k}(\tau),\mathbf{y}(\tau))$ constructed with the diffusion denoise model $\epsilon_{\theta}$. Meanwhile, the trainable forward and backward diffusion models $\epsilon_{\theta}$ in our BiTrajDiff framework are parameterized with $\theta^{b}$ and $\theta^{f}$ respectively. Thus, given the condition $\mathbf{y}(\tau)$, the forward and backward diffusion models are able to generate the denoised trajectory $\mathbf{x}_{0}(\tau)$. This process begins with sampling random noise $\mathbf{x}_{K}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, followed by iterative denoising steps where $\mathbf{x}_{k-1}\sim p_{\theta}$, for $k=\{K,\cdots,1\}$.
To realize conditional diffusion model training and inference, we take the choice of Classifier-Free Guidance (CFG) technique [18], which uses only a single denoise model to handle both conditional and unconditional generation. Formally, the perturbed noise after CFG can be represented as:
$\hat{\epsilon}=\omega*\epsilon_{\theta}(\mathbf{x}_{k}(\tau),k,\mathbf{y}(\tau))+(1-\omega)*\epsilon_{\theta}(\mathbf{x}_{k}(\tau),k,\emptyset),$ (3)
where the scalar $\omega\in[0,1]$ denotes the guidance weight of CFG. Setting $\omega$ to 00 degrades the generation to unconditional generation. Conversely, an increased value of $\omega$ promotes a stronger adherence of the generated samples to the provided condition $\mathbf{y}(\tau)$. To train such diffusion denoise models with CFG, the diffusion denoisers are trained with an objective that enables it to predict noise conditioned on $\mathbf{y}(\tau)$ and also unconditionally. This is typically achieved by randomly dropping the condition $\mathbf{y}(\tau)$ during training with probability $p$. Specifically, we have the following loss:
$\mathcal{L}_{\text{denoise}}(\theta)=\mathbb{E}_{\tau\sim\mathcal{D},k\sim\mathcal{U}\{1,K\},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\beta\sim\text{Bern}(p)}\left[\parallel\!\epsilon-\epsilon_{\theta}(\mathbf{x}_{k}(\tau),k,(1-\beta)\mathbf{y}(\tau)+\beta\emptyset)\!\parallel^{2}\right],$ (4)
where $\text{Bern}(p)$ represents the Bernoulli distribution. In our implementation, we fix the state condition $s_{t}$ in $\mathbf{x}_{k}(\tau)$ during the adding-noise training as well as the denoising inference stage for simplicity and efficiency following [1, 21, 7, 47], while the diffusion denoise models $\epsilon_{\theta}$ only take the cumulative return $R(\tau)$ components of $\mathbf{y}(\tau)$ as condition inputs.
4.2 Bidirectional Trajectory Generation
Based on the pretrained bidirectional diffusion model, we introduce a novel data-augmentation pipeline through three key stages: bidirectional state trajectory generation, bidirectional trajectory completion, and bidirectional trajectory filtering. The pipeline synthesizes global trajectories, thereby establishing connections between states not accessible in the original datasets. This process extends behavioral coverage while upholding physical consistency, ultimately producing augmented data of requisite quality for integration with offline RL algorithms.
4.2.1 Bidirectional State Trajectory Generation
We utilize the pretrained forward and backward trajectory diffusion model in section 4.1 to generate bidirectional state trajectories. Prior methods [30, 27] model only the conditional distribution of trajectories originating from a given state, neglecting the historical pathways leading to those states. As a result, these data augmentation techniques yield only limited expansion in behavior diversity. To address this limitation without introducing additional noise, our proposed method, BiTrajDiff, generates both forward and backward trajectories and integrates them by stitching at a shared intermediate state $s_{t}$, which serves as a consistent anchor.
Specifically, we sample the candidate state $s$ from the offline dataset $\mathcal{D}$ as the condition state $s_{t}$. By combining the state $s_{t}$ with the manually selected return signal $R(\tau)$ into the $\mathbf{y}(\tau)$, we can leverage the pre-trained bidirectional diffusion model $\theta$ to generate two distinct trajectories: a forward-future trajectory $\mathbf{x}^{f}_{0}(\tau)$ originates from $s_{t}$, and a backward-history trajectory $\mathbf{x}^{b}_{0}(\tau)$ concludes at $s_{t}$. Since $s_{t}$ is both the start state of $\mathbf{x}^{f}_{0}(\tau)$ and the end state of $\mathbf{x}^{b}_{0}(\tau)$, we can directly stitch them to construct the bidirectional state trajectory $\hat{\tau}_{s}(s_{t})$. Formally, let $\mathbf{x}_{0}(\tau)[i:j]$ represents the observation slice from time $i$ to time $j$, the stitched bidirectional state trajectory $\hat{\tau}_{s}(s_{t})$ can be formulated as:
$\hat{\tau}_{s}(s_{t})$ =$=\left\{\mathbf{x}^{b}_{0}(\tau)[0:H-1],s_{t},\mathbf{x}^{f}_{0}(\tau)[1:H]\right\}$ (5)
=$=\left\{\hat{s}_{t-H+1},\cdots,\hat{s}_{t-1},s_{t},\hat{s}_{t+1},\cdots,\hat{s}_{t+H-1}\right\},$ (6)
where $H$ is the horizon, and $\hat{s}_{i}$ is the diffusion generated states at MDP timestep $i$. By conditioning on and concatenating at the intermediate anchor state $s_{t}$, we construct a state trajectory, $\hat{\tau}_{s}(s_{t})$, that connects $\hat{s}_{t-H+1}$ and $\hat{s}_{t+H-1}$. This process enables the connection of states that may be disconnected in the original dataset, thereby substantially enriching the diversity of observed behavioral patterns.
4.2.2 Bidirectional Trajectory Completion
In this completion process, the generated bidirectional state trajectories $\hat{\tau}_{s}(s_{t})$ are further completed by adding the action and reward signal. We introduce the Inverse Dynamics Model (IDM): $f_{\psi}(s_{t},s_{t+1}):\mathcal{S}\times\mathcal{S}\rightarrow\mathcal{A}$ and the Reward Model (RM): $r_{\phi}(s_{t},a_{t}):\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. The supervised training objective of the two models is:
$\mathcal{L}(\psi,\phi)=\mathbb{E}_{(s_{t},a_{t},s_{t+1})\sim\mathcal{D}}\left[\parallel\!f_{\psi}(s_{t},s_{t+1})-a_{t}\!\parallel^{2}+\parallel\!r_{\phi}(s_{t},a_{t})-r_{t}\!\parallel^{2}\right]$ (7)
Thus, by feeding the generated state trajectory $\hat{\tau}_{s}(s_{t})$ to the trained IDM and RM sequentially, we obtain the completed bidirectional trajectory: $\hat{\tau}(s_{t})\!=\!\left\{(\hat{s}_{i},\hat{a}_{i},\hat{r}_{i},\hat{s}_{i+1})\right\}_{i=t-H+1}^{t+H-2}$ [^1] For the sake of clearity, the condition state $s_{t}$ is writen as $\hat{s}_{t}$., where $\hat{a}_{i}=f_{\psi}(\hat{s}_{i},\hat{s}_{i+1})$ and $\hat{r}_{i}=r_{\phi}(\hat{s}_{i},\hat{a}_{i})$. Therefore, we can obtain a generated dataset $\mathcal{\hat{D}}=\{\hat{\tau}(s_{t_{i}})\}_{i=1}^{\hat{N}}$, where $\hat{N}$ is the number of generate trajectories.
4.2.3 Bidirectional Trajectory Filter
To ensure dataset quality, we employ a two-stage filtering mechanism that removes both out-of-distribution and suboptimal trajectories. This selective process preserves the reliability of the augmented data without sacrificing diversity. In BiTrajDiff, the trajectory filter comprises an OOD trajectory filter and a greedy trajectory filter.
OOD Trajectory Filter
We model the OOD trajectory filter as a novelty detection problem [35], which addresses the identification of data instances that exhibit a significant deviation from the offline training dataset. And we utilize the Isolation Forest model [28] for OOD transition detection. Isolation forest assigns an anomaly score to each data instance, reflecting its susceptibility to isolation with shorter path lengths in randomly partitioned trees. Specifically, we construct the isolation forest on the original offline dataset $\mathcal{D}$, which is able to give each observation $\hat{s}_{i}$ an anomaly score $d(\hat{s}_{i})$. We defined the OOD score of a generated trajectory $d(\hat{\tau})=\sum_{i=t-H+1}^{t+H-2}d(\hat{s}_{i})$. Then we sort $\mathcal{\hat{D}}$ by $d(\hat{\tau})$ and retain the top-$C_{\text{ood}}$ trajectories with the smallest values, where $C_{\text{ood}}$ is a hyperparamter.
Greedy Trajectory Filter
The greedy trajectory filter selects the generated trajectories with the highest cumulative reward. Specifically, we sort the $C_{\text{ood}}$ trajectories retained from OOD trajectory filter by the sum reward $R(\bar{\tau})=\sum_{i=t-H+1}^{t+H-2}\hat{r}_{i}$. The top-$C_{\text{greedy}}$ trajectories with the highest sum rewards are picked as the final generated trajectories dataset $\tilde{\mathcal{D}}$ for offline RL training.
In summary, our BiTrajDiff operates in two key stages. Initially, it trains a bidirectional diffusion model to capture the distributions of both forward-future and backward-history trajectories. Subsequently, this trained model generates bidirectional trajectory sequences, all conditioned on intermediate anchor states. These generated sequences undergo further completion and filtering to create a high-fidelity augmented dataset. This augmented data is then integrated with the original datasets to train downstream offline RL algorithms. In our implementation, we iteratively sample batches of intermediate anchor states from the offline dataset $\mathcal{D}$, and generate bidirectional trajectories conditioned on these states using the aforementioned pipeline. The resulting trajectories are then incorporated into the final augmented dataset $\tilde{\mathcal{D}}$. We adopt CleanDiffuser [7] as the base backbone for BiTrajDiff, owing to its generality and robustness.
5 Experiments
To demonstrate the effectiveness of the proposed BiTrajDiff method, we conduct experiments on the D4RL benchmark [9]. Our evaluation seeks to answer the following questions: (1) Does BiTrajDiff outperform existing DA methods across various offline RL baselines? (Section 5.2) (2) Can BiTrajDiff generate more diverse trajectories compared to single-direction DA approaches? (Section 5.3)
5.1 Experiment Settings
Tasks and Datasets We conduct experiments on the D4RL benchmark [9], focusing on the Gym locomotion tasks. Specifically, we evaluate the performance of BiTrajDiff on three environments in MuJoCo tasks: HalfCheetah, Hopper, and Walker2d. Each environment includes three dataset types: medium, medium-replay, and medium-expert. The medium and medium-replay are collected from interactions of an unconverged policy with the environment, while the medium-expert dataset comprises a mixture of expert demonstrations and suboptimal trajectories.
Baselines and Evaluation Algorithms We compare our method with three state-of-the-art diffusion-based DA methods: Synther [30],DiffStitch [26] and RTDiff [47]. Meanwhile, to evaluate the effectiveness of the augmented trajectories of each DA method, we conduct extensive experiments utilizing two representative offline RL algorithms as base learners: IQL [23] and TD3+BC [10]. Both of the methods have been widely adopted as standard baselines in recent offline RL literature [36] due to their stable performance.
Table 1: Test returns of our proposed BiTrajDiff and baselines on the Gym tasks. $\pm$ corresponds to the standard deviation of the performance on 5 random seeds. The best and the second-best results of each setting are marked as bold and underline, respectively.
Source Task  IQL [23]  TD3BC [10]
  Base RTDiff Synther DiffStitch Ours  Base RTDiff Synther DiffStitch Ours
medium halfcheetah    48.2
$\pm$0.2
48.3
$\pm$0.1
49.6 $\pm$0.3 49.4
$\pm$0.14
48.6
$\pm$0.2
48.4
$\pm$0.2
49.6
$\pm$0.2
49.9
$\pm$1.2
50.4 $\pm$0.5 50.3
$\pm$0.4
hopper    67.0
$\pm$1.9
78.8
$\pm$6.0
72.0
$\pm$4.5
71.0
$\pm$4.2
81.1 $\pm$5.0   60.4
$\pm$3.7
69.7
$\pm$3.3
63.4
$\pm$4.2
60.3
$\pm$4.9
79.0 $\pm$3.5
walker2d    77.6
$\pm$5.3
81.2
$\pm$5.3
84.7
$\pm$5.5
83.2
$\pm$2.2
86.7 $\pm$1.7   82.0
$\pm$2.7
85.7
$\pm$2.3
84.8
$\pm$1.4
83.4
$\pm$1.7
86.7 $\pm$1.0
medium expert halfcheetah    90.1
$\pm$4.4
95.3 $\pm$0.2 93.3
$\pm$2.6
94.4
$\pm$1.4
95.3 $\pm$0.1   92.3
$\pm$5.1
91.4
$\pm$2.0
87.2
$\pm$11.1
96.0
$\pm$0.5
96.3 $\pm$0.4
hopper    90.4
$\pm$12.8
111.2 $\pm$3.8 90.8
$\pm$17.9
110.9
$\pm$2.9
110.9
$\pm$0.6
94.2
$\pm$11.1
108.1
$\pm$4.3
105.4
$\pm$9.7
107.1
$\pm$7.0
109.0 $\pm$5.1
walker2d    112.4
$\pm$0.4
111.7
$\pm$0.3
111.4
$\pm$0.7
111.6
$\pm$0.1
112.5 $\pm$0.5   109.3
$\pm$0.2
110.2 $\pm$0.1 110.2 $\pm$0.5 110.2 $\pm$0.3 110.2 $\pm$0.3
medium replay halfcheetah    43.5
$\pm$0.2
45.2
$\pm$0.2
46.6 $\pm$0.2 44.7
$\pm$0.1
44.1
$\pm$0.2
44.1
$\pm$0.1
49.0 $\pm$0.5 45.9
$\pm$0.9
44.7
$\pm$0.3
45.0
$\pm$0.6
hopper    94.7
$\pm$4.0
107.4 $\pm$1.7 103.2
$\pm$0.4
102.1
$\pm$0.4
102.8
$\pm$0.6
64.4
$\pm$8.9
71.8
$\pm$15.2
53.4
$\pm$15.5
79.6
$\pm$13.5
85.0 $\pm$10.3
walker2d    73.3
$\pm$7.5
83.4
$\pm$2.5
83.3
$\pm$5.9
86.6
$\pm$2.8
87.6 $\pm$2.2   80.9
$\pm$6.0
87.9
$\pm$3.1
90.5 $\pm$4.3 89.7
$\pm$4.2
89.9
$\pm$1.4
Average  77.5 84.7 83.8 84.1 85.5  75.1 75.8 80.2 80.2 83.5
5.2 Main Results
Results for Gym Tasks. The experimental results on D4RL Gym tasks are summarized in Table 1. Our proposed BiTrajDiff demonstrates consistent superiority across all dataset configurations. Specifically, it achieves notable performance gains in all of the environments under both TD3BC and IQL algorithms, while simultaneously maintaining high stability. Moreover, BiTrajDiff consistently outperforms existing DA methods across most scenarios. In terms of final average performance, BiTrajDiff achieves gains of 10.3% on IQL and 11.2% on TD3BC, substantially exceeding improvements reported by alternative methods like DiffStitch. These consistent gains across both IQL and TD3BC backbones underscore the effectiveness and generality of our BiTrajDiff approach.
5.3 Visulization
Figure 1: Visualization of the comparative dynamic error and L2 distance metrics for BiTrajDiff-generated trajectories versus forward and backward diffusion-generated trajectories, all evaluated with a consistent horizon.
To evaluate the accuracy and diversity robustness of the BiTrajDiff framework, we compare its generated trajectories with those produced by single-direction forward and backward diffusion models. Following [30], we employ two quantitative metrics to assess a generated trajectory $\hat{\tau}$: 1) Dynamic Error: $\mathcal{E}_{\text{Dyn}}(\hat{\tau})=\sum_{t}\|\hat{s}_{t+1}-s_{t+1}\|_{2}$, which measures trajectory accuracy by summing the mean squared error between the predicted next states $\hat{s}_{t+1}$ and the ground truth $s_{t+1}$; 2) L2 Distance: $\mathcal{E}_{\text{L2D}}(\hat{\tau})=\min_{\tau\in\mathcal{D}}\sum_{t}\|\hat{s}_{t}-s_{t}\|_{2}$, defined as the minimal L2 distance between the generated trajectory and any trajectory of equal length in the dataset $\mathcal{D}$, capturing trajectory diversity. As shown in Figure 1, while forward and backward diffusion models achieve low $\mathcal{E}_{\text{Dyn}}$, they exhibit limited $\mathcal{E}_{\text{L2D}}$, indicating that the generated trajectories closely resemble those in the original dataset and thus offer limited performance gains. In contrast, trajectories generated by BiTrajDiff demonstrate significantly broader marginal distributions in $\mathcal{E}_{\text{L2D}}$ while maintaining competitive $\mathcal{E}_{\text{Dyn}}$, reflecting enhanced behavioral diversity without sacrificing dynamic consistency. Consequently, BiTrajDiff facilitates greater performance gains with the augmented data. These results highlight the ability of BiTrajDiff to generate behaviorally diverse trajectories while preserving accuracy, effectively balancing exploration and reliability in offline trajectory generation.
6 Conclusion
In this paper, we present BiTrajDiff, a novel diffusion-based framework that enhances offline RL by enabling bidirectional trajectory generation from shared intermediate states. Unlike previous single-direction augmentation methods that only model forward future trajectories from observed states, BiTrajDiff simultaneously generates both forward and backward trajectories. This dual-generation strategy allows BiTrajDiff to not only imagine plausible futures but also reconstruct potential historical transitions, thereby creating novel trajectories that globally connect states, including those unreachable in the original dataset. This leads to significantly diverse and effective trajectory augmentation, improving offline RL performance beyond the capability of the forward-only generation schemes of other DA frameworks. Extensive experiments on the D4RL benchmark also demonstrate that BiTrajDiff significantly outperforms existing data augmentation baselines across diverse offline RL algorithms. Future work will explore bidirectional generation for offline RL under multi-task settings with broader behavioral patterns.